OpenStudy (anonymous):

How do you integrate 3cos^2*5xdx?

OpenStudy (vishweshshrimali5):

\[\large{\int 3\cos^2 (?) * 5x dx}\] What is the angle of cos^2 here ?

OpenStudy (anonymous):

@vishweshshrimali5 I think he means \[\int\limits_{}^{} 3( \cos ^{2} (5x))dx\]

OpenStudy (vishweshshrimali5):

Possible :) Then: \[\large{I = \int 3\cos^2 (5x) \ dx}\tag{1}\] First of all substitute: \[\large{u = 5x \implies du = 5dx}\tag{2}\]

OpenStudy (vishweshshrimali5):

We will get: \[\large{I = \cfrac{3}{5}\int \cos^2 u \ du}\]

OpenStudy (anonymous):

Watching and learning as you're doing it.

OpenStudy (vishweshshrimali5):

Use this: \[\large{\cos (2\theta) = 2\cos^2 {\theta} - 1}\] \[\large{\implies 2\cos^2 \theta = \cos(2\theta) + 1}\] \[\large{\implies \cos^2 \theta = \cfrac{\cos(2\theta) + 1}{2}}\]

OpenStudy (anonymous):

oh double angle identity?

OpenStudy (vishweshshrimali5):

Now using this we get: \[\large{I = \cfrac{3}{5} \int \cfrac{\cos (2u) + 1}{2} \ du}\] \[\large{\implies I = \cfrac{3}{10} \int [\cos (2u) + 1] \ du}\] @study100 Yes :)

OpenStudy (anonymous):

Ok so far I'm understanding this :)) <3 why abs value?

OpenStudy (anonymous):

wait, that sign is not abs. Sorry kinda blind.

OpenStudy (vishweshshrimali5):

Now we can simplify I as: \[\large{I = \cfrac{3}{10}\left[\int \cos(2u) \ du + \int du \right]}\] \[\large{\implies I = \cfrac{3}{10}\left[\cfrac{\sin (2u)}{2} + u\right] + C}\]

OpenStudy (anonymous):

you're amazing. : o

OpenStudy (vishweshshrimali5):

Now we replace u by 5x: \[\large{\boxed{\color{green}{I = \cfrac{3}{10}\left[\cfrac{\sin (10x)}{2} + 5x\right] + C}}}\]

OpenStudy (vishweshshrimali5):

@study100 Thanks :)

OpenStudy (vishweshshrimali5):

So using a little bit substitution and trigonometric identities we can solve this question.

OpenStudy (anonymous):

I wish I can medal this like 20 times.

OpenStudy (vishweshshrimali5):

:) Thanks

OpenStudy (anonymous):

Thanks a lot, this really helped